OpenStudy (anonymous):

How would i answer the equation : x+1/3x-6 = 5x/6 + 1/ x-2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x = 4 - 3·√2 and x = 3·√2 + 4

OpenStudy (anonymous):

so do i set the equation equal to zero first before i start using the quadratic equation?

OpenStudy (anonymous):

yeah!

OpenStudy (radar):

Subtract 1/(x-2) from both sides getting: x+1 1 5x ---- - ----- = ----- 3(x-2) (x-2) 6 Perform subtraction on the left using LCD of 3(x-2) x+1 3 5x ----- - ------ = ---- 3(x-2) 3(x-2) 6 combine left hand side and reduce fractions x-2 5x ---- = -- 3(x-2) 6 1/3=5x/6 cross multiply 15x=6 x=6/15=2/5